I jave restarted my math... My problem is = 3x+2 = 6x+4
Answer by Theo(13342)

2x+2 = 6x+4
subtract 2x from both sides of the equation to get 2x - 2x + 2 = 6x + 4 - 2x
simplify by combining like terms to get 2 = 4x + 4
subtract 4 from both sides of the equation to get 2 - 4 = 4x + 4 - 4
simplify by combining like terms to get -2 = 4x
divide both sides of the equation by 4 to get -2/4 = 4x/4
simplify by performing the operations indicated to get -1/2 = x
your solution is x = -1/2
replace x in your original equation with -1/2 and the equation should be true.
start with 2x + 2 = 6x + 4
replace x with -1/2 to get 2*(-1/2) + 2 = 6*(-1/2) + 4
simplify by performing the operations indicated to get -2/2 + 2 = -6/2 + 4
simplify further by completing the operations indicated to get -1 + 2 = -3 + 4
combine like terms to get 1 = 1.
since 1 = 1 is true, your solution of x = -1/2 is good.